# Kronecker-product random matrices and a matrix least squares problem

### v1

@@ 4.2. Weak estimates in operator norm
Note that $\Im(\mathsf{z}-i(\delta/2)+z)>\delta/2$ for all $z\in\mathbb{C}^{+}$, so by Lemma D.1, this inverse is well-defined and
@@
We now check that $\mathbb{P}[\mathcal{E}]\geq 1-n^{-D}$ for all $n\geq n_{0}(l,K,D)$: By a standard tail bound for the operator norm (see e.g. [23, Theorem 7.3]), the event
@@
where we have used also $\|R_{ii}^{(S)}\|_{\mathrm{op}}\leq\|R^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ by Lemma D.1 and the assumption $\|\mathsf{x}\|_{\mathrm{op}}\leq\gamma$. Adjusting the value of $\alpha$ yields the second statement of (77).
@@
On $\mathcal{E}$, we have $\|\mathrm{I}\|_{\mathrm{op}}\leq n^{-\alpha}$ by (68). For $\mathrm{II}$ and $\mathrm{IV}$, Lemma D.1 implies $\|\widetilde{R}_{ii}^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ for both $i\in S$ and $i\notin S$, the latter because $\|R_{ii}^{(S)}\|_{\mathrm{op}}\leq\|R^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ and the former because $\Im R^{(S)}\geq 0$ so $\Im R_{jj}^{(S)}\geq 0$ and $\Im\mathsf{z}_{i}\geq\delta$ in the definition (79). Then
@@ 4.3. Iterative bootstrapping
Applying this, the identity $\mathsf{a}^{-1}-\mathsf{b}^{-1}=\mathsf{a}^{-1}(\mathsf{b}-\mathsf{a})\mathsf{b}^{-1}$, and $\|\widehat{R}_{ii}\|_{\mathrm{op}},\|(-\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x})^{-1}\|_{\mathrm{op}}\leq\delta^{-1}$ by Lemma D.1, we may approximate $\widehat{R}_{ii}$ in (104) as
@@ 6. Analysis of least-squares problem
For any $\mathsf{d}_{1},\ldots,\mathsf{d}_{k+1}\in\mathcal{D}$, by the free moment-cumulant relations [51, Proposition 11.4], we have
@@
where $\operatorname{NC}(2k+1)$ is the set of non-crossing partitions of $(1,2,\ldots,2k+1)$ and $\kappa_{\pi}$ is the free cumulant associated to each $\pi\in\operatorname{NC}(2k+1)$. Since $\mathsf{a}$ is free of $\mathcal{D}$ and has $2^{\text{nd}}$ free cumulant equal to 1 and remaining free cumulants 0 [51, Example 11.21], we have $\kappa_{\pi}(\mathsf{d}_{1},\mathsf{a},\mathsf{d}_{2},\mathsf{a},\ldots,\mathsf{d}_{k},\mathsf{a},\mathsf{d}_{k+1})=0$ unless the blocks of $\pi$ containing $\mathsf{a}$ constitute a non-crossing pairing and are disjoint from those containing $(\mathsf{d}_{1},\ldots,\mathsf{d}_{k+1})$. Let $\operatorname{NC}_{2}(k)$ denote the set of all non-crossing pairings of $\{2,4,6,\ldots,2k\}$ corresponding to the locations of the $\mathsf{a}$’s, and for each $\rho\in\operatorname{NC}_{2}(k)$, let $K(\rho)\in\operatorname{NC}(k+1)$ be its complement in $\{1,2,\ldots,2k+1\}$, i.e. the coarsest non-crossing partition of the remaining elements $\{1,3,5,\ldots,2k+1\}$ for which $\rho\cup K(\rho)$ forms a non-crossing partition of $\{1,2,\ldots,2k+1\}$. Then each $\pi\in\operatorname{NC}(2k+1)$ for which $\kappa_{\pi}(\mathsf{d}_{1},\mathsf{a},\mathsf{d}_{2},\mathsf{a},\ldots,\mathsf{d}_{k},\mathsf{a},\mathsf{d}_{k+1})\neq 0$ is the union of some $\rho\in\operatorname{NC}_{2}(k)$ and some $\bar{\rho}\leq K(\rho)$ that refines $K(\rho)$, so we have
@@ Appendix B Operator concentration inequalities
Let $\mathcal{Y}\subset\mathcal{X}$ be a von Neumann subalgebra with $\phi$-invariant conditional expectation $\phi^{\mathcal{Y}}:\mathcal{X}\to\mathcal{Y}$. Suppose $\mathsf{x}_{1},\ldots,\mathsf{x}_{n}\in L^{p}(\mathcal{X})$ satisfy $\phi^{\mathcal{Y}}(\mathsf{x}_{i})=0$, and are independent over $\mathcal{Y}$ in the sense that for each $i$, each $\mathsf{x}$ in the von Neumann subalgebra generated by $\mathsf{x}_{i}$, and each $\mathsf{x}^{\prime}$ in the von Neumann subalgebra generated by $\{\mathsf{x}_{j}:j\neq i\}$, we have $\phi^{\mathcal{Y}}(\mathsf{x}\mathsf{x}^{\prime})=\phi^{\mathcal{Y}}(\mathsf{x})\phi^{\mathcal{Y}}(\mathsf{x}^{\prime})$.
@@
By [14, Theorem 1], for a universal constant $C>0$,
@@
For (a), fix any $p\geq 2$. We apply Lemma B.1 in the setting of [38, Example 1.3]: Let $L^{\infty}(\Omega)$ be the von Neumann algebra of bounded scalar random variables over the underlying probability space $(\Omega,\mathscr{F},\mathbb{P})$, and consider $\mathcal{M}=L^{\infty}(\Omega)\otimes\mathcal{X}$ equipped with the state $\mathbb{E}\circ\phi$. Then $\alpha_{i}\mathsf{x}_{i}\in L^{p}(\mathcal{M},\mathbb{E}\circ\phi)$, $\mathbb{E}:\mathcal{M}\to\mathcal{X}$ coincides with the conditional expectation onto the subalgebra $\mathcal{X}\subset\mathcal{M}$, and $\{\alpha_{i}\mathsf{x}_{i}\}_{i=1}^{n}$ are independent over $\mathcal{X}$ in the sense of Lemma B.1, so Lemma B.1 shows
@@
Here, for $p\geq 2$, the third term is bounded by the second by the following argument (see also [38, Eq. (2.4)]): Consider $\hat{\mathsf{x}}=\sum_{i=1}^{n}E_{i1}\otimes\mathsf{x}_{i}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$ equipped with the trace $n^{-1}\operatorname{Tr}\otimes\phi$, and the linear map $T:\mathbb{C}^{n\times n}\otimes\mathcal{X}\to\mathbb{C}^{n\times n}\otimes\mathcal{X}$ defined by $T(E_{ij}\otimes\mathsf{x})=E_{i,j+i-1}\otimes\mathsf{x}$ where $j+i-1$ is interpreted modulo $n$. Thus $\hat{\mathsf{x}}$ has $\mathcal{X}$-valued entries $\mathsf{x}_{1},\ldots,\mathsf{x}_{n}$ along the first column, and $T(\hat{\mathsf{x}})$ has these entries instead along the main diagonal. We have
@@ Appendix C Fluctuation averaging
In this section, we prove Lemma 3.5. The proof is analogous to the argument in the scalar setting of [29, Lemma A.2]. For part (c), we will apply the following combinatorial lemma from [29].
@@
We collect here several pieces of background on von Neumann algebras and non-commutative $L^{p}$ spaces that are needed in our main arguments. We refer to [49], [13, Section 1], and [57, Chapter 14] for additional discussion. Throughout, $\mathcal{X}$ is a (finite) von Neumann algebra with faithful, normal, tracial state $\phi$ as in Section 3.
Lemma D.1 ([34], Lemma 3.1).
Suppose $\mathsf{x}\in\mathcal{X}$ is such that for some $\epsilon>0$, either $\Im\mathsf{x}\geq\epsilon 1$ or $\Im\mathsf{x}\leq-\epsilon 1$. Then $\mathsf{x}$ is invertible and $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq 1/\epsilon$.
@@
Defining $\|\mathsf{x}\|_{p}=\phi(|x|^{p})^{1/p}$, the space $L^{p}(\mathcal{X})$ is the Banach space completion of $\mathcal{X}$ under $\|\cdot\|_{p}$. We set $\|\mathsf{x}\|_{\infty}\equiv\|\mathsf{x}\|_{\mathrm{op}}$ and $L^{\infty}(\mathcal{X})\equiv\mathcal{X}$. These spaces $L^{p}(\mathcal{X})$ may be continuously embedded into a common space of (unbounded) densely-defined operators affiliated to $\mathcal{X}$ — we refer to [49] or [13, Section 1] for this construction.
@@
For $p=1$, let $\mathsf{y}\in\mathcal{B}$ be the unitary operator for which $\mathsf{y}\phi^{\mathcal{B}}(\mathsf{x})=|\phi^{\mathcal{B}}(\mathsf{x})|$. Then

### v2

@@ 4.2. Weak estimates in operator norm
Note that $\Im(\mathsf{z}-i(\delta/2)+z)>\delta/2$ for all $z\in\mathbb{C}^{+}$, so by Lemma 3.6, this inverse is well-defined and
@@
We now check that $\mathbb{P}[\mathcal{E}]\geq 1-n^{-D}$ for all $n\geq n_{0}(l,K,D)$: By a standard tail bound for the operator norm (see e.g. [24, Theorem 7.3]), the event
@@
where we have used also $\|R_{ii}^{(S)}\|_{\mathrm{op}}\leq\|R^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ by Lemma 3.6 and the assumption $\|\mathsf{x}\|_{\mathrm{op}}\leq\gamma$. Adjusting the value of $\alpha$ yields the second statement of (77).
@@
On $\mathcal{E}$, we have $\|\mathrm{I}\|_{\mathrm{op}}\leq n^{-\alpha}$ by (68). For $\mathrm{II}$ and $\mathrm{IV}$, Lemma 3.6 implies $\|\widetilde{R}_{ii}^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ for both $i\in S$ and $i\notin S$, the latter because $\|R_{ii}^{(S)}\|_{\mathrm{op}}\leq\|R^{(S)}\|_{\mathrm{op}}\leq\delta^{-1}$ and the former because $\Im R^{(S)}\geq 0$ so $\Im R_{jj}^{(S)}\geq 0$ and $\Im\mathsf{z}_{i}\geq\delta$ in the definition (79). Then
@@ 4.3. Iterative bootstrapping
Applying this, the identity $\mathsf{a}^{-1}-\mathsf{b}^{-1}=\mathsf{a}^{-1}(\mathsf{b}-\mathsf{a})\mathsf{b}^{-1}$, and $\|\widehat{R}_{ii}\|_{\mathrm{op}},\|(-\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x})^{-1}\|_{\mathrm{op}}\leq\delta^{-1}$ by Lemma 3.6, we may approximate $\widehat{R}_{ii}$ in (104) as
@@ 6. Analysis of least-squares problem
For any $\mathsf{d}_{1},\ldots,\mathsf{d}_{k+1}\in\mathcal{D}$, by the free moment-cumulant relations [52, Proposition 11.4], we have
@@
where $\operatorname{NC}(2k+1)$ is the set of non-crossing partitions of $(1,2,\ldots,2k+1)$ and $\kappa_{\pi}$ is the free cumulant associated to each $\pi\in\operatorname{NC}(2k+1)$. Since $\mathsf{a}$ is free of $\mathcal{D}$ and has $2^{\text{nd}}$ free cumulant equal to 1 and remaining free cumulants 0 [52, Example 11.21], we have $\kappa_{\pi}(\mathsf{d}_{1},\mathsf{a},\mathsf{d}_{2},\mathsf{a},\ldots,\mathsf{d}_{k},\mathsf{a},\mathsf{d}_{k+1})=0$ unless the blocks of $\pi$ containing $\mathsf{a}$ constitute a non-crossing pairing and are disjoint from those containing $(\mathsf{d}_{1},\ldots,\mathsf{d}_{k+1})$. Let $\operatorname{NC}_{2}(k)$ denote the set of all non-crossing pairings of $\{2,4,6,\ldots,2k\}$ corresponding to the locations of the $\mathsf{a}$’s, and for each $\rho\in\operatorname{NC}_{2}(k)$, let $K(\rho)\in\operatorname{NC}(k+1)$ be its complement in $\{1,2,\ldots,2k+1\}$, i.e. the coarsest non-crossing partition of the remaining elements $\{1,3,5,\ldots,2k+1\}$ for which $\rho\cup K(\rho)$ forms a non-crossing partition of $\{1,2,\ldots,2k+1\}$. Then each $\pi\in\operatorname{NC}(2k+1)$ for which $\kappa_{\pi}(\mathsf{d}_{1},\mathsf{a},\mathsf{d}_{2},\mathsf{a},\ldots,\mathsf{d}_{k},\mathsf{a},\mathsf{d}_{k+1})\neq 0$ is the union of some $\rho\in\operatorname{NC}_{2}(k)$ and some $\bar{\rho}\leq K(\rho)$ that refines $K(\rho)$, so we have
@@ Appendix B Operator concentration inequalities
Let $\mathcal{Y}\subset\mathcal{X}$ be a von Neumann subalgebra with $\phi$-invariant conditional expectation (c.f. Lemma D.1) $\phi^{\mathcal{Y}}:\mathcal{X}\to\mathcal{Y}$. Suppose $\mathsf{x}_{1},\ldots,\mathsf{x}_{n}\in L^{p}(\mathcal{X})$ satisfy $\phi^{\mathcal{Y}}(\mathsf{x}_{i})=0$, and are independent over $\mathcal{Y}$ in the sense that for each $i$, each $\mathsf{x}$ in the von Neumann subalgebra generated by $\mathsf{x}_{i}$, and each $\mathsf{x}^{\prime}$ in the von Neumann subalgebra generated by $\{\mathsf{x}_{j}:j\neq i\}$, we have $\phi^{\mathcal{Y}}(\mathsf{x}\mathsf{x}^{\prime})=\phi^{\mathcal{Y}}(\mathsf{x})\phi^{\mathcal{Y}}(\mathsf{x}^{\prime})$.
@@
By [15, Theorem 1], for a universal constant $C>0$,
@@
For (a), fix any $p\geq 2$. We apply Lemma B.1 in the setting of [39, Example 1.3]: Let $L^{\infty}(\Omega)$ be the von Neumann algebra of bounded scalar random variables over the underlying probability space $(\Omega,\mathscr{F},\mathbb{P})$, and consider $\mathcal{M}=L^{\infty}(\Omega)\otimes\mathcal{X}$ equipped with the state $\mathbb{E}\circ\phi$. Then $\alpha_{i}\mathsf{x}_{i}\in L^{p}(\mathcal{M},\mathbb{E}\circ\phi)$, $\mathbb{E}:\mathcal{M}\to\mathcal{X}$ coincides with the conditional expectation onto the subalgebra $\mathcal{X}\subset\mathcal{M}$, and $\{\alpha_{i}\mathsf{x}_{i}\}_{i=1}^{n}$ are independent over $\mathcal{X}$ in the sense of Lemma B.1, so Lemma B.1 shows
@@
Here, for $p\geq 2$, the third term is bounded by the second by the following argument (see also [39, Eq. (2.4)]): Consider $\hat{\mathsf{x}}=\sum_{i=1}^{n}E_{i1}\otimes\mathsf{x}_{i}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$ equipped with the trace $n^{-1}\operatorname{Tr}\otimes\phi$, and the linear map $T:\mathbb{C}^{n\times n}\otimes\mathcal{X}\to\mathbb{C}^{n\times n}\otimes\mathcal{X}$ defined by
$T(E_{ij}\otimes\mathsf{x})=\begin{cases}E_{i,j+i-1}\otimes\mathsf{x}&\text{if $j=1$}\\ 0&\text{otherwise}\end{cases}$
where $j+i-1$ is interpreted modulo $n$. Thus $\hat{\mathsf{x}}$ has $\mathcal{X}$-valued entries $\mathsf{x}_{1},\ldots,\mathsf{x}_{n}$ along the first column, and $T(\hat{\mathsf{x}})$ has these entries instead along the main diagonal. We have
@@ Appendix C Fluctuation averaging
In this section, we prove Lemma 3.5. The proof is analogous to the argument in the scalar setting of [30, Lemma A.2]. For part (c), we will apply the following combinatorial lemma from [30].
@@
We collect here several pieces of background on von Neumann algebras and non-commutative $L^{p}$ spaces that are needed in our main arguments. We refer to [50], [14, Section 1], and [58, Chapter 14] for additional discussion. Throughout, $\mathcal{X}$ is a (finite) von Neumann algebra with faithful, normal, tracial state $\phi$ as in Section 3.
@@
Defining $\|\mathsf{x}\|_{p}=\phi(|x|^{p})^{1/p}$, the space $L^{p}(\mathcal{X})$ is the Banach space completion of $\mathcal{X}$ under $\|\cdot\|_{p}$. We set $\|\mathsf{x}\|_{\infty}\equiv\|\mathsf{x}\|_{\mathrm{op}}$ and $L^{\infty}(\mathcal{X})\equiv\mathcal{X}$. These spaces $L^{p}(\mathcal{X})$ may be continuously embedded into a common space of (unbounded) densely-defined operators affiliated to $\mathcal{X}$ — we refer to [50] or [14, Section 1] for this construction.
@@
For $p=1$, let $\mathsf{y}\in\mathcal{B}$ be the projection operator for which $\mathsf{y}\phi^{\mathcal{B}}(\mathsf{x})=|\phi^{\mathcal{B}}(\mathsf{x})|$. Then